OpenStudy (anonymous):

If y varies inversely with x and y = 3 when x = 4, what is the value of y when x = 2? A.1/6 B.y = 2 C.y = 6 D.y = 12

OpenStudy (anonymous):

@jdoe0001 can you help me again?

OpenStudy (radar):

y = kx this is a direct variation, k being a constant of proportion y = k/x this is an inverse variation. k still being the constant of proportion do you see the difference?

OpenStudy (anonymous):

Yeah, the direct variation is the multiply and the inverse is the divide i think?

OpenStudy (radar):

Yes. The first step is solve for the k. You are given that 3 = k/4 can you solve for k?

OpenStudy (anonymous):

0.75 if i divide. 12 if i multiply

OpenStudy (radar):

What divided by 4 equals 3? That is what the equation is saying.

OpenStudy (radar):

One of your answers is correct.

OpenStudy (anonymous):

i multiply both side by 4 i get 12.

OpenStudy (radar):

Yes, k is 12. Now we can solve the problem y = 12/2

OpenStudy (anonymous):

How to do that?

OpenStudy (anonymous):

if i divide 12/2 i get 6y

OpenStudy (radar):

They have given you the value of x as 2, and you now know k=12 so it is quite easy to solve for y.